OpenStudy (tayaiv):

Two dice are tossed together. What is the probability of getting an odd number on the second die and a total of five?

OpenStudy (tayaiv):

@welshfella Could you help me, please?

OpenStudy (shelbywyatt):

3/12?

OpenStudy (tayaiv):

@ShelbyWyatt That's not one of the options. The options are: 1/8 11/18 7/18 4/9

OpenStudy (tayaiv):

@FaiqRaees Could you maybe help me with this, please?

OpenStudy (shelbywyatt):

my guess is 1/8

OpenStudy (mrnood):

a guess is not much use please show some working

OpenStudy (shelbywyatt):

⇒ odds in favour of E1 = P(E1)/[1 – P(E1)] = (1/9)/(1 – 1/9) = 1/8.

OpenStudy (mrnood):

It seems to me that the only throws that produce the result are 2,3 4,1 which would resuult in 2/36 as the odds but that is not an option - so maybe 3,2 1,4 also count giving 4/36 Hmm - not getting this...

OpenStudy (shelbywyatt):

section i

OpenStudy (welshfella):

you could list all the favourable results first second 2 3 4 1 which gives a probability of 2 / 36

OpenStudy (welshfella):

the second die must give an odd number

OpenStudy (mrnood):

the dice are thrown 'together' so strictly there is no 'second die so as I said above 4/36 seems an option - but it is not in the answer....

OpenStudy (welshfella):

AH yes together . so you can count any one of them as a second die. Correct.

OpenStudy (shelbywyatt):

Let E1 be the event of getting the sum 5. Then, E1 = {(1, 4), (2, 3), (3, 2), (4, 1)} ⇒ P(E1) = 4 Therefore, P(E1) = n(E1)/n(S) = 4/36 = 1/9 ⇒ odds in favour of E1 = P(E1)/[1 – P(E1)] = (1/9)/(1 – 1/9) = 1/8.

OpenStudy (shelbywyatt):

any of you understand, cuz i dont

OpenStudy (welshfella):

hmmm But the question asks for the probability

OpenStudy (mrnood):

it doesnt ask for the 'odds' it asks for the probability I understand your point - and it matcjes an answer - but not the way the question is asked in my mind. E.g. the probability of throwing a 6 on 1 die is 1:6 The ODDS of throwing a 6 is 1:5

OpenStudy (shelbywyatt):

yeah..

OpenStudy (welshfella):

yae the odds are given as a ratio

OpenStudy (welshfella):

Are you sure you have the correct options, @tayaiv ?

OpenStudy (tayaiv):

@welshfella

OpenStudy (shelbywyatt):

... -.-

OpenStudy (tayaiv):

Thank you all for trying to figure this out with me, by the way. I really appreciate your help!

OpenStudy (shelbywyatt):

so you would have to guess it?

OpenStudy (tayaiv):

@ShelbyWyatt No, I'm still trying to figure it out myself. I won't guess

OpenStudy (tayaiv):

I think the answer is 1/18. @mrnood was right: 4/36 includes all of the possibilities of getting 5, if we consider (1,4) and (4,1) different options. The probability of getting an odd number on the second die is 3/6 or 1/2. So, 1/2 * 4/36 = 1/18 What do you think?

OpenStudy (shelbywyatt):

sound good

OpenStudy (tayaiv):

Thank you all!!

OpenStudy (mrnood):

I don't agree- the 2 results are not independent you cannot throw a 5 on 1 dice and get a sum of 5

OpenStudy (mrnood):

and 1/18 is not an answer option anyhow

OpenStudy (welshfella):

I think the question has not been worded very well. If it said a die is tossed twice the answer would definitely be 1/18. I would put 1/18 as the answer anyway.

OpenStudy (mrnood):

I still disagree the calculation is based on a 0.5 proability of thraowin ODD on second die. BUT a throw of 5 is NOT a success (because the toatal cannot be 5) The calculation used above is that for independent events (i.e. what is the chance of throwin 2 dice and scoring 5 AND throwing 1 dice and getting ODD)

OpenStudy (shelbywyatt):

lol

OpenStudy (welshfella):

Surely its a TOTAL of 5 - the favourable results are 1 and 4 , and 2 and 3 which is 2 results out of a possible 36.

OpenStudy (mrnood):

the scoring of 5 total on 2 dice thrown together has a probability of 1/9 (1,4) (4,1) (2,3) (3,2) HOWEVER this cannot be combined with the probability of thraowin ODD on one of the dice (0.5) because they are not independent - i.e. you cannot throw a 5 on 1 dice AND score a total of 5. Therefore 1/9* 1/2 is not vallid because this is only valid for independent events. (and 1/18 is not an option anyhow) @tayaiv are your answers correct (you wrote 11/18 - is it possible that should be 1/18??)

OpenStudy (welshfella):

1/18 is an option given in the picture that tayaiv posted